# Reduction-induced Variation of Partial Von Neumann Entropy

### v2

@@ IV Its Superiorities in comparison to existing bipartite QE measures
which represents the upper limit of QE as derived from the mathematical structure of the system’s density matrix and reflect that they may not accurately reflect the degree of mixed-state QE, and the Negativity expressed as
$\operatorname{N}(\rho)=\frac{\sqrt{3+\cos{4\alpha}}-1-\cos{2\alpha}}{2}.$ (25)
The five aforementioned measures are illustrated in Fig.2. Notably, when $\alpha=\frac{\pi}{2}$, the following relations hold: $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. For values $\alpha=0.8,2.341$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ hold; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, we find $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, the relations become $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$. Clearly, for the specified mixed two-qubit quantum systems, the monotonic behavior of these five measures is consistent. Specifically, the differences in the values of Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence, and Negativity arise from their distinct physical implications and mathematical structures. Regarding the bipartite QE degree metric for the specified mixed states, it is concluded that the Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation, and Negativity exhibit advantages over the Relative Entropy of Entanglement and Concurrence. When considering bipartite QE degree metric in mixed states, the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy, is superior to Entanglement of Formation, Negativity, Relative Entropy of Entanglement and Concurrence.
Fig. 2: Reduction-induced Variation of Partial Von Neumann entropy as well as other QE measures such as Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$. For values $\alpha=0.8,2.341$, $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ ; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$.
@@ V Conclusion
In this paper, we propose a method called Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in bipartite mixed states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states especially for any mixed states. This method exhibits minimal computational complexity, making it highly efficient. Moreover, it has a wide range of applicability. Its intuitive and clear physical representation, combined with simple computation and wide range of applications, indicates the ease of exploring its specific potential applications. By considering the degree of bipartite QE, along with the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy proves to be a good measure for assessing bipartite mixed-state QE compared to established methods. In future studies, the conceptual framework of the Reduction-Induced Variation of Partial Von Neumann Entropy should be further developed to describe and quantify multipartite QE.

### v4

@@ IV Its Superiorities in comparison to existing bipartite QE measures
which represents the upper limit of the bipartite mixed-state QE as derived from the mathematical structure of the system’s density matrix. The aforementioned four kinds of bipartite QE measures are depicted in Fig.2. Clearly, for the specified bipartite quantum systems, the monotonic behavior of these measures is consistent. Notably, when $\alpha=\frac{\pi}{2}$, the following relations hold: $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. While within the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, the relations become $\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$. The Relative Entropy of Entanglement and Concurrence, representing the upper limit of QE, are the largest among these measures, and reflect that they may not accurately reflect the degree of bipartite mixed-state QE. Conversely, the Reduction-induced Variation of Partial Von Neumann Entropy is the smallest, with the Entanglement of Formation falling in between. These differences in metric values, arise from their distinct physical implications and mathematical structures. Regarding the bipartite QE degree metric for mixed states, it is concluded that the Reduction-induced Variation of Partial Von Neumann Entropy and Entanglement of Formation exhibit advantages over the Relative Entropy of Entanglement and Concurrence. When considering bipartite QE degree metric in mixed states, the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy is superior to the Entanglement of Formation, Relative Entropy of Entanglement and Concurrence.
Fig. 2: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as Entanglement of Formation, Relative Entropy of Entanglement and Concurrence varying with $\alpha$. When $\alpha=\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. While within the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, $\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$.
@@ V Conclusion
In this paper, we propose a method called Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite sates, especially for mixed states. Partial Von Neumann Entropy is merely a special case of this method, serving as a measure of bipartite pure-state QE. This method exhibits minimal computational complexity, making it highly efficient. Moreover, it has a wide range of applicability. Its intuitive and clear physical representation, along with simple computation and wide range of applications, indicates the ease of exploring its specific potential applications. By considering the degree of bipartite QE, along with the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy proves to be a good measure for assessing bipartite mixed-state QE compared to other established methods like Entanglement of Formation, Relative Entropy of Entanglement and Concurrence. In future studies, the conceptual framework and line of thought behind the Reduction-Induced Variation of Partial Von Neumann Entropy deserves to be further developed to describe and quantify multipartite QE.